OpenStudy (anonymous):

5 (x • y) = (5 • x) y is an example of which algebraic property? Associative Property of Multiplication Transitive Property Reflexive Property Commutative Property of Multiplication

OpenStudy (anonymous):

Associative Property of Multiplication Because \[(A \times B) \times C = A \times (B \times C)\]

OpenStudy (anonymous):

They can associate with anyone

OpenStudy (anonymous):

ya associate because the parentheses just changed.